OpenStudy (anonymous):

Evaluate the following expression for n = -4 and p = -3 -5n – 4p + 7n + 2

OpenStudy (vishweshshrimali5):

I can help you in this

OpenStudy (vishweshshrimali5):

If n = -4 and p = -3 Then -5n - 4p + 7n + 2 = -5(-4) - 4(-3) + 7(-4) + 2 = 20 + 12 - 28 + 2 = 6

OpenStudy (vishweshshrimali5):

Ok ARE YOU SATISFIED WITH THE ANSWER

OpenStudy (anonymous):

yes thank you

OpenStudy (vishweshshrimali5):

Please become my fan And add my question as good answer Will you ? Please

OpenStudy (anonymous):

Evaluate the following expression for c = -7 and d = 5 2d + 5c – 3d + 2dc

OpenStudy (anonymous):

help me with thhat please